OpenStudy (anonymous):

Write the equation in slope intercept form of the line the has a slope of 2 and contains the point (1,1)

OpenStudy (anonymous):

First we write the general form of the equation for a line i.e. y=mx + b where m = slope ( which is given as 2) So equation modifies to y=2x+b. To find b, plug in (1,1) for x and y Now you get 1=2(1)+b. 1 = 2 + b 1 - 2 = b So b=-1. So the final equation is y=2x-1 So y = 2x - 1 is the line with slope 2 and containing the point (1,1)